# Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs

### v2

@@ 5 Related Work
The theoretical beginnings of citation analysis can be traced back to foundational works such as [9] identification of reasons for citation [23] studies on citation function. Early annotation schemes, such as those by [29], were later adapted by [32] for supervised machine learning approaches to citation classification.
[17] introduced the ACL-ARC dataset, which contains nearly citations from papers in the NLP field, annotated for their function with a classification scheme of six classes. [25] extended this classification scheme by refining the comparison class to capture similarities, differences, and disagreement. Around the same time, [5] proposed a multitask model incorporating structural information from scientific papers. They also introduced the SciCite dataset, which is significantly larger and spans multiple scientific domains, with three intent classes.
[3] introduced SciBERT, a BERT-based [7] encoder language model pre-trained specifically on scientific text, which has since become the backbone of many citation intent classification methods. SciBERT has been widely adopted due to its ability to generalize across scientific domains. For example, [22] introduced ImpactCite, an XLNet-based method for citation impact analysis, which was later used by [24] to achieve state-of-the-art results on the SciCite dataset. Paolini et al. demonstrated the effectiveness of ensemble classifiers combining fine-tuned SciBERT and XLNet models.
Recent research has continued to explore PLM-based methods for citation intent classification. [20] used a prompt-based learning approach on SciBERT to identify citation intent, while [28] achieved state-of-the-art performance on the ACL-ARC dataset by proposing a multi-task learning framework that jointly fine-tunes SciBERT on a dataset of primary interest together with multiple auxiliary datasets to take advantage of additional supervision signals. [19] explored various prompting and tuning strategies on SciBERT, including fixed and dynamic context prompts, and found that parameter updating with prompts improved performance. They also briefly experiment on LLMs by evaluating the zero-shot performance of GPT-3.5, which performed well on their recently introduced ACT2 dataset but poorly on the ACL-ARC dataset. However, GPT-3.5 was not evaluated on SciCite. In contrast, our work is the first to focus entirely on evaluating and fine-tuning numerous open-weight, general-purpose LLMs without using models pre-trained explicitly and exclusively for the scientific domain.
@@
Method | SciCite | ACL-ARC |
Feature-rich Random Forest [17] | – | 53.00 |
Structural Scaffolds [5] | 84.00 | 67.90 |
SciBERT [3] | 85.22 | – |
ImpactCite [22] | 88.93 | – |
CitePrompt [20] | 86.33 | 68.39 |
EnsIntWS [24] | 89.46 | – |
EnsIntWoS [24] | 88.48 | – |
MTL Finetuning (Search) [28] | 85.25 | 64.56 |
MTL Finetuning (TRL) [28] | 85.35 | 75.57 |

### v3

@@ 4 Related Work
The theoretical beginnings of citation analysis can be traced back to foundational works such as Garfield et al.’s (1965) identification of reasons for citation and Moravcsik and Murugesan’s (1975) studies on citation function. Early annotation schemes, such as those by Spiegel-Rosing (1977), were later adapted by Teufel et al. (2006) for supervised machine learning approaches to citation classification.
Jurgens et al. (2018) introduced the ACL-ARC dataset, which contains nearly citations from papers in the NLP field, annotated for their function with a classification scheme of six classes. Pride and Knoth (2020) extended this classification scheme by refining the comparison class to capture similarities, differences, and disagreement. Around the same time, Cohan et al. (2019) proposed a multitask model incorporating structural information from scientific papers. They also introduced the SciCite dataset, which is significantly larger and spans multiple scientific domains, with three intent classes.
Method | SciCite | ACL-ARC |
Feature-rich RF (Jurgens et al., 2018) | – | 53.00 |
Structural Scaffolds (Cohan et al., 2019) | 84.00 | 67.90 |
SciBERT (Beltagy et al., 2019) | 85.22 | – |
ImpactCite (Mercier et al., 2021) | 88.93 | – |
CitePrompt (Lahiri et al., 2023) | 86.33 | 68.39 |
EnsIntWS (Paolini et al., 2024) | 89.46 | – |
EnsIntWoS (Paolini et al., 2024) | 88.48 | – |
MTL Finetuning (Search) (Shui et al., 2024) | 85.25 | 64.56 |
MTL Finetuning (TRL) (Shui et al., 2024) | 85.35 | 75.57 |
Beltagy et al. (2019) introduced SciBERT, a BERT-based (Devlin et al., 2019) encoder language model pre-trained specifically on scientific text, which has since become the backbone of many citation intent classification methods. SciBERT has been widely adopted due to its ability to generalize across scientific domains. For example, Mercier et al. (2021) introduced ImpactCite, an XLNet-based method for citation impact analysis, which was later used by Paolini et al. (2024) to achieve state-of-the-art results on the SciCite dataset. Paolini et al. demonstrated the effectiveness of ensemble classifiers combining fine-tuned SciBERT and XLNet models.
Recent research has continued to explore PLM-based methods for citation intent classification. Lahiri et al. (2023) used a prompt-based learning approach on SciBERT to identify citation intent, while Shui et al. (2024) achieved state-of-the-art performance on the ACL-ARC dataset by proposing a multi-task learning framework that jointly fine-tunes SciBERT on a dataset of primary interest together with multiple auxiliary datasets to take advantage of additional supervision signals. Kunnath et al. (2023) explored various prompting and tuning strategies on SciBERT, including fixed and dynamic context prompts, and found that parameter updating with prompts improved performance. They also briefly experiment on LLMs by evaluating the zero-shot performance of GPT-3.5, which performed well on their recently introduced ACT2 dataset but poorly on the ACL-ARC dataset. However, GPT-3.5 was not evaluated on SciCite. In contrast, our work is the first to focus entirely on evaluating and fine-tuning numerous open-weight, general-purpose LLMs without using models pre-trained explicitly and exclusively for the scientific domain.